OpenStudy (anonymous):

solve equation by square root property 3(x+7)^2=60

OpenStudy (anonymous):

browneez, divide through by 3 first to get\[(x+7)^2=20\]Then take the square root of both sides \[\sqrt{(x+7)^2}=\pm \sqrt{20} \rightarrow x+7=\pm \sqrt{20} \rightarrow x=-7 \pm \sqrt{20}\]Now, you have that \[20=4 \times 5 = 2^2 \times 5\]so that\[\sqrt{20}=\sqrt{2^2 \times 5}=\sqrt{2^2}\times \sqrt{5}=2\sqrt{5}\]so \[x=-7 \pm 2\sqrt{5}\]

OpenStudy (anonymous):

You have to take a 'plus or minus' on one side when you take the square root. We usually attach it to the number.

OpenStudy (anonymous):

thanks, the 3 on the outside of the() was throwing me off.